# Low-depth quantum symmetrization

### v1

@@ Appendix A An exact SIL symmetrization algorithm
Our exact SIL symmetrization algorithm relies on a classical parallel algorithm for generating a random permutation proposed by Alonso and Schott (1996). However, that algorithm was introduced in a non-reversible way. Therefore, in this appendix, we present that algorithm in a pedagogical manner and indicate along the way how to make it reversible.
A.1 Lower exceeding sequences
@@
The central tool used in the Alonso and Schott (1996) algorithm is the concept of lower exceeding sequences (LESs).
@@
The number of LESs of length $n$ is $n!$, while the number of permutations of $n$ items is also $|S_{n}|=n!$. Reference Alonso and Schott (1996) efficiently constructs a bijective mapping between these sets and uses it to generate a random permutation by generating an LES and then applying the mapping. Our quantum algorithm is based on a similar approach.
The mapping of Alonso and Schott (1996) is defined as follows.
Definition 4 ($s^{\sigma}$ as in Alonso and Schott (1996)).
@@
It is easy to verify that $s^{\sigma}$ is an LES for any $\sigma\in S_{n}$, but it is less intuitive to explain the meaning of the values in $s^{\sigma}$. Fortunately, Alonso and Schott (1996) give an intuitive way to understand the meaning of $s^{\sigma}$ diagrammatically, as follows.
@@
A naive implementation of the above procedure for converting an LES to its corresponding permutation takes $O(n\log n)$ depth. Fortunately, Alonso and Schott (1996) provides a $\operatorname{poly}(\log n)$-depth algorithm to do the conversion, which is outlined next.
A.2 From LES to permutation with reversible merging
Alonso and Schott (1996) show that the blocks can be filled in a divide-and-conquer manner, as follows. Divide the LES into two parts at the middle, make the sub-diagrams for both parts individually, and merge them together. The first two steps are straightforward, but the “merging” step is nontrivial. A simple merging strategy is not hard to see: one can insert every row of the left diagram into the unfilled rows of the right diagram. More specifically, we first concatenate two diagrams together and keep the right diagram unchanged. Then, if the $j$th column of the left diagram has the $y_{j}$th block filled, we move this block to the $y_{j}$th unfilled row of the right diagram.
@@ A.2 From LES to permutation with reversible merging
Performing this merging naively takes $O(n)$ steps for each pair of sub-diagrams. To reduce it to $O(\log n)$ steps, Alonso and Schott (1996) introduce another way of representing the diagrams: a sorted list of $(r_{i},c_{i})$ tuples, where each $(r_{i},c_{i})$ represents the row and column coordinates of a filled block in the diagram. We describe the general algorithm together with a running example that shows how it works in the case depicted in Fig. 5.
@@
The key strategy of Alonso and Schott (1996) is to first calculate the $\Gamma_{f}$-sorted tuple lists for the left and right sub-diagrams, and merge them together to build the whole diagram. Note that each sub-diagram’s list can also be calculated recursively using the same divide-and-conquer method. Without loss of generality, we only describe the last step of the algorithm, i.e., merging the last two sub-diagrams to obtain $T_{f}$, since all previous steps use the same procedure. We also emphasize that we describe the whole algorithm in a reversible manner—unlike in the original paper Alonso and Schott (1996)—so that it can be directly quantized.
@@
$T_{1}=(r_{\omega(1)},c_{\omega(1)})(r_{\omega(2)},c_{\omega(2)})\dots(r_{\omega(2k)},c_{\omega(2k)})=(r_{\omega(1)},z_{1})(r_{\omega(2)},z_{2})\dots(r_{\omega(2k)},z_{2k}).$ (81)
@@
Step 1: $T_{0}\mapsto T_{1}$. We reversibly sort $T_{0}$ to obtain $T_{1}$ using comparison-based sorting algorithms, based on the rules given by Alonso and Schott (1996). First, we define $\tilde{T}_{0}$ and $\tilde{T}_{1}$ by concatenating each tuple in the lists with its index:
@@
$\tilde{T}_{1}$ $=(\omega(1),r_{\omega(1)},c_{\omega(1)})(\omega(2),r_{\omega(2)},c_{\omega(2)})\dots(\omega(2k),r_{\omega(2k)},c_{\omega(2k)})$
$=(\omega(1),r_{\omega(1)},z_{1})(\omega(2),r_{\omega(2)},z_{2})\dots(\omega(2k),r_{\omega(2k)},z_{2k}).$
@@
According to Ref. Alonso and Schott (1996), $\tilde{T}_{1}$ is a $\Gamma_{1}$-sorted tuple list, where $\Gamma_{1}$ is defined as follows. (Here we simply describe $\Gamma_{1}$ without proving why $\tilde{T}_{1}$ is $\Gamma_{1}$-sorted; see Ref. Alonso and Schott (1996) for the proof.) For $\omega(i)\neq\omega(j)$, $(\omega(i),r_{\omega(i)},c_{\omega(i)})<_{\Gamma_{1}}(\omega(j),r_{\omega(j)},c_{\omega(j)})$ if and only if any of the following conditions is satisfied:
@@
One can now get $\tilde{T}_{1}$ by sorting $\tilde{T}_{0}$ using any reversible sorting network without knowing $\omega$ explicitly, since $\tilde{T}_{1}=\mathcal{A}_{\Gamma_{1}}(\tilde{T}_{0})$ and $\tilde{T}_{0}=\mathcal{A}_{\Gamma_{0}}(\tilde{T}_{1})$. This step takes $O(\log n)$ depth.
@@
A.3 The SIL symmetrization algorithm
Since we can implement the merging process ($T_{0}\mapsto T_{1}\mapsto T_{f}$) in depth $O(\log^{2}n)$ ($O(\log n)$ for $\tilde{T}_{0}\mapsto\tilde{T}_{1}$, $O(\log^{2}n)$ for $\tilde{T}_{1}\mapsto T_{1}$, and $O(\log^{2}n)$ for $T_{1}\mapsto T_{f}$), and this divide-and-conquer algorithm uses $O(\log n)$ layers of merging, the total depth of the LES-to-permutation conversion is $O(\log^{3}n)$. To generate the superposition of all permutations of the nontrivial input $l\in\mathrm{SIL}$ with $|l|=n$ and $l\neq(123\dots n)$ in low depth, we “move” the permutations stored in $\sigma(12\dots n)$ to $l$ using the following quantum algorithm. Note that all ancilla qubits involved are omitted.
@@ A.3 The SIL symmetrization algorithm
    Generate permutations from LESs using the quantized Alonso and Schott (1996) algorithm:
@@
    $\frac{1}{\sqrt{n!}}\ket{00\dots 0}_{1}\sum_{\sigma\in S_{n}}\ket{\sigma(12\dots n)}_{2}\stackrel{{\scriptstyle\mathrm{SORT}_{2,1}}}{{\longrightarrow}}\frac{1}{\sqrt{n!}}\sum_{\sigma\in S_{n}}\ket{\operatorname{rec}(\sigma)}_{1}\ket{12\dots n}_{2}$ (87)
@@
    $\frac{1}{\sqrt{n!}}\sum_{\sigma\in S_{n}}\ket{\operatorname{rec}(\sigma)}_{1}\ket{l}_{2}\stackrel{{\scriptstyle\mathrm{UNSORT}_{2,1}}}{{\longrightarrow}}$ $\frac{1}{\sqrt{n!}}\sum_{\sigma\in S_{n}}\ket{00\dots 0}_{1}\ket{\sigma(l)}_{2}$ (88)
@@
Appendix B The exact NSIL symmetrization algorithm
The only difference between the exact algorithm (Result 2) and the $O(\log n)$-depth algorithm (Result 1) is how they prepare the equal superposition over the $H_{l}$ subgroup. Therefore, in this appendix, we focus on this task. Our strategy consists of two steps: detecting duplicates, and preparing random permutations of duplicates in superposition.
B.1 Duplicate detection algorithm
We now describe an $O(\log n)$-depth quantum algorithm to detect all duplicated items in an ordered list. This algorithm takes as input |$\ket{l}_{1}\ket{0}_{2}$ where $l$ is an NSIL and each integer in $l$ is assigned $O(\log n)$ qubits in register 2 to record how many times this integer is repeated in $l$. The output of this algorithm is |$\ket{l}\ket{\operatorname{dup}(l)}$ where $\operatorname{dup}(l)$ records all repetitions: if there exists $n_{i}>1$ for $i$ such that $l_{i-1}<l_{i}=l_{i+1}=\dots=l_{i+n_{i}-1}<l_{i+n_{i}}$, then $\operatorname{dup}(l)_{i}=n_{i}$; for all other cases, $\operatorname{dup}(l)_{i}=0$. For example, if $l=(13333338)$, then $\operatorname{dup}(l)=(06000000)$.
@@ B.1 Duplicate detection algorithm
B.2 Subgroup superposition preparation
@@
Each $h\in H_{l}$ can be generated from an $s\in\mathrm{LES}_{l}$, and each $s\in\mathrm{LES}_{l}$ can be computed from an $h\in H_{l}$, using the rules in Appendix A.
@@
One can use the rules in Appendix A to calculate the corresponding $s^{h}\in\mathrm{LES}_{l}$ from an $h\in H_{l}$ as follows. Since permutations only exist between repeated elements, and repeated elements are adjacent in $l\in\mathrm{NSIL}_{n}$, we observe that
@@
Therefore, $s^{h}\in\mathrm{LES}_{l}$. This immediately implies that one can use the rules in Appendix A to find the corresponding $h\in H_{l}$ from $s^{h}$. ∎

### v2

@@ IV.1 An exact SIL symmetrization algorithm
Our exact SIL symmetrization algorithm relies on a classical parallel algorithm for generating a random permutation proposed by [13]. However, that algorithm was introduced in a non-reversible way. Therefore, in this subsection, we present that algorithm in a pedagogical manner and indicate along the way how to make it reversible.
Lower exceeding sequences
@@
The central tool used in the Alonso and Schott [13] algorithm is the concept of lower exceeding sequences (LESs).
@@
The number of LESs of length $n$ is $n!$, while the number of permutations of $n$ items is also $|S_{n}|=n!$. Reference [13] efficiently constructs a bijective mapping between these sets and uses it to generate a random permutation by generating an LES and then applying the mapping. Our quantum algorithm is based on a similar approach.
The mapping of [13] is defined as follows.
Definition 4 ($s^{\sigma}$ as in [13]).
@@
It is easy to verify that $s^{\sigma}$ is an LES for any $\sigma\in S_{n}$, but it is less intuitive to explain the meaning of the values in $s^{\sigma}$. Fortunately, Alonso and Schott [13] give an intuitive way to understand the meaning of $s^{\sigma}$ diagrammatically, as follows.
@@
A naive implementation of the above procedure for converting an LES to its corresponding permutation takes $O(n\log n)$ depth. Fortunately, [13] provides a $\operatorname{poly}(\log n)$-depth algorithm to do the conversion, which is outlined next.
From LES to permutation with reversible merging
Alonso and Schott [13] show that the blocks can be filled in a divide-and-conquer manner, as follows. Divide the LES into two parts at the middle, make the sub-diagrams for both parts individually, and merge them together. The first two steps are straightforward, but the “merging” step is nontrivial. A simple merging strategy is not hard to see: one can insert every row of the left diagram into the unfilled rows of the right diagram. More specifically, we first concatenate two diagrams together and keep the right diagram unchanged. Then, if the $j$th column of the left diagram has the $y_{j}$th block filled, we move this block to the $y_{j}$th unfilled row of the right diagram.
@@
Performing this merging naively takes $O(n)$ steps for each pair of sub-diagrams. To reduce it to $O(\log n)$ steps, Alonso and Schott [13] introduce another way of representing the diagrams: a sorted list of $(r_{i},c_{i})$ tuples, where each $(r_{i},c_{i})$ represents the row and column coordinates of a filled block in the diagram. We describe the general algorithm together with a running example that shows how it works in the case depicted in Fig. 4.
@@
The key strategy of Alonso and Schott [13] is to first calculate the $\Gamma_{f}$-sorted tuple lists for the left and right sub-diagrams, and merge them together to build the whole diagram. Note that each sub-diagram’s list can also be calculated recursively using the same divide-and-conquer method. Without loss of generality, we only describe the last step of the algorithm, i.e., merging the last two sub-diagrams to obtain $T_{f}$, since all previous steps use the same procedure. We also emphasize that we describe the whole algorithm in a reversible manner—unlike in the original paper [13]—so that it can be directly quantized.
@@
$T_{1}=$ $(r_{\omega(1)},c_{\omega(1)})(r_{\omega(2)},c_{\omega(2)})\dots(r_{\omega(2k)},c_{\omega(2k)})$ (49)
$=$ $(r_{\omega(1)},z_{1})(r_{\omega(2)},z_{2})\dots(r_{\omega(2k)},z_{2k}).$
@@
Step 1: $T_{0}\mapsto T_{1}$. We reversibly sort $T_{0}$ to obtain $T_{1}$ using comparison-based sorting algorithms, based on the rules given by Alonso and Schott [13]. First, we define $\tilde{T}_{0}$ and $\tilde{T}_{1}$ by concatenating each tuple in the lists with its index:
@@
$\tilde{T}_{1}$ $=(\omega(1),r_{\omega(1)},c_{\omega(1)})(\omega(2),r_{\omega(2)},c_{\omega(2)})\dots$
$\quad\dots(\omega(2k),r_{\omega(2k)},c_{\omega(2k)})$
$=(\omega(1),r_{\omega(1)},z_{1})(\omega(2),r_{\omega(2)},z_{2})\dots$
$\quad\dots(\omega(2k),r_{\omega(2k)},z_{2k}).$
@@
According to Ref. [13], $\tilde{T}_{1}$ is a $\Gamma_{1}$-sorted tuple list, where $\Gamma_{1}$ is defined as follows. (Here we simply describe $\Gamma_{1}$ without proving why $\tilde{T}_{1}$ is $\Gamma_{1}$-sorted; see Ref. [13] for the proof.) For $\omega(i)\neq\omega(j)$, $(\omega(i),r_{\omega(i)},c_{\omega(i)})<_{\Gamma_{1}}(\omega(j),r_{\omega(j)},c_{\omega(j)})$ if and only if any of the following conditions is satisfied:
@@
One can now get $\tilde{T}_{1}$ by sorting $\tilde{T}_{0}$ using any reversible sorting network without knowing $\omega$ explicitly, since $\tilde{T}_{1}=\mathcal{A}_{\Gamma_{1}}(\tilde{T}_{0})$ and $\tilde{T}_{0}=\mathcal{A}_{\Gamma_{0}}(\tilde{T}_{1})$. This step takes $\tilde{O}(\log n)$ depth.
@@
The SIL symmetrization algorithm
Since we can implement the merging process ($T_{0}\mapsto T_{1}\mapsto T_{f}$) in depth $\tilde{O}(\log^{2}n)$ ($\tilde{O}(\log n)$ for $\tilde{T}_{0}\mapsto\tilde{T}_{1}$, $\tilde{O}(\log^{2}n)$ for $\tilde{T}_{1}\mapsto T_{1}$, and $\tilde{O}(\log^{2}n)$ for $T_{1}\mapsto T_{f}$), and this divide-and-conquer algorithm uses $O(\log n)$ layers of merging, the total depth of the LES-to-permutation conversion is $\tilde{O}(\log^{3}n)$. To generate the superposition of all permutations of the nontrivial input $l\in\mathrm{SIL}$ with $|l|=n$ and $l\neq(123\dots n)$ in low depth, we “move” the permutations stored in $\sigma(12\dots n)$ to $l$ using the following quantum algorithm. Note that all ancilla qubits involved are omitted.
@@
    Generate permutations from LESs using the quantized Alonso and Schott [13] algorithm:
@@
    $\frac{1}{\sqrt{n!}}\ket{00\dots 0}_{1}\sum_{\sigma\in S_{n}}\ket{\sigma(12\dots n)}_{2}$ (55)
    $\stackrel{{\scriptstyle\mathrm{SORT}_{2,1}}}{{\longrightarrow}}$ $\frac{1}{\sqrt{n!}}\sum_{\sigma\in S_{n}}\ket{\operatorname{rec}(\sigma)}_{1}\ket{12\dots n}_{2}$
@@
    $\frac{1}{\sqrt{n!}}\sum_{\sigma\in S_{n}}\ket{\operatorname{rec}(\sigma)}_{1}\ket{l}_{2}$ (56)
    $\stackrel{{\scriptstyle\mathrm{UNSORT}_{2,1}}}{{\longrightarrow}}$ $\frac{1}{\sqrt{n!}}\sum_{\sigma\in S_{n}}\ket{00\dots 0}_{1}\ket{\sigma(l)}_{2}$
@@
IV.2 NSIL symmetrization for superposed inputs
Recall that our algorithm for NSIL symmetrization involves preparing the equal superposition over the $H_{l}$ subgroup. Our strategy consists of two steps: detecting duplicates and preparing random permutations of duplicates in superposition.
@@ IV.2 NSIL symmetrization for superposed inputs
Duplicate detection algorithm
We now describe an $\tilde{O}(\log n)$-depth quantum algorithm to detect all duplicated items in an ordered list. This algorithm takes as input |$\ket{l}_{1}\ket{0}_{2}$ where $l$ is an NSIL and each integer in $l$ is assigned $O(\log n)$ qubits in register 2 to record how many times this integer is repeated in $l$. The output of this algorithm is |$\ket{l}\ket{\operatorname{dup}(l)}$ where $\operatorname{dup}(l)$ records all repetitions: if there exists $n_{i}>1$ for $i$ such that $l_{i-1}<l_{i}=l_{i+1}=\dots=l_{i+n_{i}-1}<l_{i+n_{i}}$, then $\operatorname{dup}(l)_{i}=n_{i}$; for all other cases, $\operatorname{dup}(l)_{i}=0$. For example, if $l=(13333338)$, then $\operatorname{dup}(l)=(06000000)$.
@@
Subgroup superposition preparation
@@
Each $h\in H_{l}$ can be generated from an $s\in\mathrm{LES}_{l}$, and each $s\in\mathrm{LES}_{l}$ can be computed from an $h\in H_{l}$, using the rules in Section IV.1.
@@
One can use the rules in Section IV.1 to calculate the corresponding $s^{h}\in\mathrm{LES}_{l}$ from an $h\in H_{l}$ as follows. Since permutations only exist between repeated elements, and repeated elements are adjacent in $l\in\mathrm{NSIL}_{n}$, we observe that
@@
Therefore, $s^{h}\in\mathrm{LES}_{l}$. This immediately implies that one can use the rules in Section IV.1 to find the corresponding $h\in H_{l}$ from $s^{h}$. ∎
@@
V Efficient conversion from second quantization to first quantization
To convert any given second-quantized state to its first-quantized counterpart, we first transform the second-quantized state to an NSIL before employing the (anti-)symmetrization algorithm on the increasing lists. To do so, we design a novel reversible circuit that takes as input a list of occupation numbers (i.e., a second-quantized representation) and outputs an increasing list of positions (or modes) occupied, where each mode is assigned with a number written in its binary representation. More specifically, our circuit transforms $W_{0}:=n_{0}n_{1}\dots n_{m-1}$ (where $n_{i}\geq 0$ is the occupation number of mode $i$) into $W_{f}:=0^{n_{0}}1^{n_{1}}\dots(m-1)^{n_{m-1}}$. Our approach is based on reversible sorting networks and takes $\tilde{O}(\log^{2}n)$ depth where $n=\sum_{i=0}^{m-1}n_{i}$. The main idea for converting $W_{0}$ to $W_{f}$ reversibly is to first generate $n$ empty registers; then use sorting networks to move $n_{i}$ of them in between of the $(i-1)$th register and the $i$th register, for all $i\in[m]$; and use the new positions of the empty registers to compute the index $i$ locally for each of them. In more detail, the algorithm proceeds as follows.
$\tilde{W}_{1}$ $=(0,n_{0},n_{0})(1,n_{1},n_{0}+n_{1})(2,n_{2},n_{0}+n_{1}+n_{2})$ (66)
$\quad\dots(m-1,n_{m-1},n).$
$\tilde{W}_{2}$ $=(0,n_{0},n_{0})(1,n_{1},n_{0}+n_{1})\dots$ (67)
$\quad(m-1,n_{m-1},n)(m,0,0)(m,0,1)\dots(m,0,n-1).$
  • $a_{1}<a_{2}$;
  • $a_{1}=a_{2}=m$ AND $c_{1}<c_{2}$.
Now, using reversible sorting networks, we sort $\tilde{W}_{2}$ to obtain
$\quad\dots(m,0,n_{0}+n_{1}-1)(1,n_{1},n_{0}+n_{1})(m,0,n_{0}+n_{1})$
$\quad\dots(m,0,n-1)(m-1,n_{m-1},n),$
  • $c_{1}<c_{2}$,
  • $c_{1}=c_{2}$ AND $a_{1}<a_{2}$.
Now, we have allocated $n_{i}$ such elements (with first element being $m$) between $(i-1,n_{i-1},\sum_{j=0}^{i-1}n_{j})$ and $(i,n_{i},\sum_{j=0}^{i}n_{j})$, and we need to assign the value $i$ to each of them. For the $k$th element in $\tilde{W}_{3}$, if the first value is $m$ (meaning that this is a newly added element) and the third value is $i_{k}$ (the index of new elements), then the corresponding $i$ value is simply $k-i_{k}$ (the difference between the index in $\tilde{W}_{3}$ and the index in all new elements). We add this value to the second value for each new element, obtaining
$\tilde{W}_{4}$ $=(m,0,0)(m,0,1)\dots(m,0,n_{0}-1)(0,n_{0},n_{0})$ (69)
$\quad(m,1,n_{0})\dots(m,1,n_{0}+n_{1}-1)(1,n_{1},n_{0}+n_{1})$
$\quad(m,2,n_{0}+n_{1})\dots(m,m-1,n-1)$
$\quad(m-1,n_{m-1},n).$
$\tilde{W}_{5}$ $=(0,n_{0},n_{0})(1,n_{1},n_{0}+n_{1})$ (70)
$\quad\dots(m-1,n_{m-1},n)(m,0,0)\dots.$
$\tilde{W}_{7}$ $=(m,0,0)\dots(m,0,n_{0}-1)(0,0,n_{0})(m,1,n_{0})\dots$ (72)
$\quad(m,1,n_{0}+n_{1}-1)(1,0,n_{0}+n_{1})(m,2,n_{0}+n_{1})$
$\quad\dots(m,m-1,n-1)(m-1,0,n).$
Next, we notice that $(i,0,\sum_{j=0}^{i}n_{i})$ must be the $(\sum_{j=0}^{i}n_{i}+i)$th element in $\tilde{W}_{7}$ for $i\in\{0,1,m-1\}$. We can therefore subtract all $\sum_{j=0}^{i}n_{i}$ values and obtain
$\tilde{W}_{8}$ $=(m,0,0)\dots(m,0,n_{0}-1)(0,0,0)(m,1,n_{0})\dots$ (73)
$\quad(m,1,n_{0}+n_{1}-1)(1,0,0)(m,2,n_{0}+n_{1})\dots$
$\quad(m,m-1,n-1)(m-1,0,0).$
Next, we can sort $\tilde{W}_{8}$ again under $\Gamma_{W_{2}}$ to obtain
$\tilde{W}_{f}$ $=(0,0,0)\dots(m-1,0,0)(m,0,0)\dots$ (74)
$\quad(m,0,n_{0}-1)(m,1,n_{0})\dots(m,1,n_{0}+n_{1}-1)\dots$
$\quad(m,m-1,n-1).$
$W_{f}=0^{n_{0}}1^{n_{1}}\dots(m-1)^{n_{m-1}}.$ (75)
Since all operations in this process are either reversible sorting or reversible local data modification, the whole process including symmetrization takes $\tilde{O}(\log^{3}n)$ depth and can be implemented as a reversible circuit.
VI Application: quantum interferometric imaging system
VI.1 QFT-based single-photon interferometry
We review quantum telescope arrays [11, 12] for the low-photon-rate scenario with the help of the schematic diagram in Fig. 7. For simplicity, we work in an idealized model of a 2-dimensional world where all $m$ telescopes are equally spaced in a line. Each telescope $T_{0},T_{1},\dots,T_{m-1}$ has a lens (or mirror) to collect photons and a well-isolated photon detector equipped with a bosonic quantum memory (since photons are bosons). The distance between adjacent telescopes is $d$. The photon source (which may be a star, galaxy, etc.) is denoted by $S$ and is assumed to be infinitely far away from the telescope array, so its position can be described using a single parameter $\theta$, the angle between $T_{0}S$ and $SP$ where $P$ is the nearest point to $S$ on the telescope plane. Since both $d$ and $m$ are finite, the following angle relation can be derived from trigonometry:
[…]
$a^{\dagger}(\theta)\ket{00\dots 0}$ (78)
[…]
We can now perform the QFT on this state to obtain
Observe that if
$\frac{d\sin\theta}{\lambda}-\frac{k}{m}=v$ (82)
$k\left(\frac{\lambda}{md}\right)=\sin\theta-\frac{v\lambda}{d}.$ (83)
Note that in the discussions of quantum telescope arrays in this paper, we assume there are no atmospheric fluctuations that may produce local phase differences. This may be reasonable if the telescope array is space-based, but for ground-based telescopes, such phases are inevitable and may affect image quality, since the phase difference between the $i$th and the $j$th telescopes is no longer strictly $2\pi d(j-i)\sin\theta/\lambda$. These fluctuations can potentially be corrected using classical methods such as adaptive optics [33], as mentioned in Ref. [34].
VI.2 Multiple-photon interferometry
  1. 1.
    the number of detectors $m$ is large;
  2. 2.
    the area of each aperture is large;
  3. 3.
    the field of view is large (which can be achieved when $d$ is very small);
  4. 4.
    the photon wavelength is long; or
  5. 5.
    the source is bright.
Now, we outline how our NSIL symmetrization algorithm makes it possible to extract all $\theta_{j}$ from |$\ket{\psi(\theta_{1},\dots,\theta_{n})}$ using only one copy of the state using Result 2. Let
$\phi_{i,j}:=\frac{2\pi dj\sin\theta_{i}}{\lambda}.$ (85)
The explicit multiple-photon state is
$\prod_{j=1}^{n}a^{\dagger}(\theta_{j})\ket{00\dots 0}$ (86)
As with the single-photon case, we convert to the first-quantized representation in order to apply the quantum Fourier transform. To do so, as described in Section V, we first transform |$\ket{n_{0}n_{1}\dots n_{m-1}}$ into a list $l$ of telescope indices in increasing order. This transformation circuit has $\tilde{O}(\log^{2}n)$ depth and uses $O(n\log m)$ ancilla qubits. As a simple example, if $m=3$, $n=4$, and $r=2101$, then the state in Equation (87) is |$\ket{121}$, which will be transformed to |$\ket{l(r)}=\ket{0112}$. Here $l(r)=0112$ is an NSIL as it is the sorted version of $r$. We also notice that the factor $\sqrt{\prod_{i=0}^{m-1}n_{i}!}$ is just $\sqrt{|H_{l(r)}|}$, square root of the order of the subgroup of $S_{n}$ preserving $l(r)$. We can thus write the transformed unnormalized state as
Next, we can apply the NSIL symmetrization algorithm in Result 2 to obtain the unnormalized first-quantized state
|$\ket{\psi_{\mathrm{S}}(\theta_{1},\dots,\theta_{n})}$ (89)
Finally, we rewrite
|$\ket{\psi_{\mathrm{S}}(\theta_{1},\dots,\theta_{n})}$ (93)
$\mathrm{QFT}^{n}\ket{\psi_{\mathrm{S}}(\theta_{1},\dots,\theta_{n})}=\frac{1}{\sqrt{n!}}\sum_{\sigma\in S_{n}}\bigotimes_{i=1}^{n}\ket{k_{\sigma(12\dots n)_{i}}}$ (94)
where $k_{\sigma(12\dots n)_{i}}=\frac{md}{\lambda}\left(\sin\theta_{\sigma(12\dots n)_{i}}-\frac{v\lambda}{d}\right)$ with $v\in\mathbb{Z}$ determined by the direction of the telescope, as in Equation (83).
It is also worth mentioning that instead of directly producing images, quantum telescope arrays can offer better accuracy in estimating specific parameters, such as the distance between binary stars, as introduced in Ref. [35]. Our multiple-photon interferometric imaging system can potentially extend the list of high-precision parameter estimations.
VII Summary & Discussion
In this paper, we studied the general quantum symmetrization problem. We identified subtleties with repetitions in the input list and pointed out issues with previous SIL symmetrization algorithms when acting on NSILs. Using sorting networks, we then proposed the first logarithmic-depth quantum algorithms for symmetrizing NSILs. These algorithms leverage the coset decomposition of the symmetric group, use an ancilla register to mediate multiplications of permutations elements, and adapt SIL symmetrization algorithms as subroutines. The single-input symmetrization algorithm features $\tilde{O}(\log n)$ depth and uses $O(n\log n+\log m)$ ancilla qubits for NSILs of $n$ integers with value up to $m$, while the superposed-input symmetrization algorithm uses $\tilde{O}(\log^{3}n)$ depth and $O(n\log n)$ ancilla qubits. We also proposed an $\tilde{O}(\log^{2}n)$-depth algorithm to transform between second-quantized states and $\mathrm{NSIL}$s, which allows for efficient conversion between second-quantized states and first-quantized states.
We distinguish between the scenario in our paper and in other algorithmic applications of symmetrization. It is well known that the graph isomorphism problem can be solved by creating a superposition over all permutations of a given graph [36]. However, our results cannot be used in that context since there is no increasing input guarantee in the setting of graph isomorphism. Indeed, if the vertices of a graph could be ordered in a canonical way, this would solve graph isomorphism.
Our algorithm requires all-to-all connectivity due to the sorting networks used, just as in Ref. [4]. A natural open problem is to find low-depth sorting networks with restricted connectivity, which could make our algorithm much more favorable for near-term devices. It might also be possible to reduce the circuit depth of the superposed-input symmetrization algorithm in Result 2 or increase the success probability of the algorithm in Result 1.
Acknowledgments